OpenStudy (anonymous):

Cara drives for 4 hours at a rate of 45 miles per hour. Later, she drives for another 3 hours so that her total distance traveled for the day is 336 miles. Find the solution set for 4(45) + 3r = 336 given the replacement set {52, 53, 54, 55} to determine the rate at which Cara traveled during the second part of her trip. (Points : 1) 52 53 54 55

OpenStudy (anonymous):

help

OpenStudy (phi):

4(45) + 3r = 336 what does 4(45) mean? any idea?

OpenStudy (anonymous):

4*45

OpenStudy (phi):

and 4*45= ?

OpenStudy (anonymous):

180

OpenStudy (phi):

so 4(45) + 3r = 336 is the same as 180 + 3r= 336 can you subtract 180 from both sides? what do you get?

OpenStudy (anonymous):

156

OpenStudy (phi):

what about the left side? in other words, write down the entire equation (or we will get lost)

OpenStudy (anonymous):

3r=156?

OpenStudy (phi):

close 180 + 3r= 336 180-180 + 3r = 336 -180 simplify both sides. 180-180 is 0 and 0+3r is 3r (adding 0 does "nothing") 3r = 156

OpenStudy (anonymous):

52?

OpenStudy (phi):

last step, divide both sides by 3. on the left side 3/3 is 1 and you get r

OpenStudy (phi):

yes r= 156/3 r= 52